Six projects have been identified for possible implementation by a company that makes dry ice blasters, machines that propel tiny dry ice pellets at supersonic speeds so they flash freeze and then lift grime, paint, rust, mold, asphalt, and other contaminants off in-place machines and a wide range of surfaces. The present worth of each project has been determined. Identify all acceptable mutually exclusive bundles if the budget limitation is $23,000. Arrange the acceptable bundles in such a way that the total PW values are in ascending order.

Project Project PW at 15%, $
L 29,000
M 11,000
N 41,000
O 35,000
P 6,000
Q 2,000
Answer #1

projects are organized in ascending order on the basis of PW as follows:

Project Project PW at 15%, $
Q 2,000
P 6,000
M 11,000
L 29,000
O 35,000
N 41,000

On the basis of PW only three projects Q, P and M are having PW less than of $23000.

So, feasible bundles are :

QPM = 2000 + 6000 + 11000 = $19000

QP = 2000+6000 = $8000

PM = 6000+11000 = $17000

So, Q, P & M projects will be selected.